# ViSTA: Visual Storytelling using Multi-modal Adapters for Text-to-Image Diffusion Models

### v2

@@ Abstract.
Figure 1. ViSTA sample visual storytelling. Given a sequence of narrative story prompts and the first text-image pair as initial references, we generate high-quality story images aligned with their corresponding prompts, maintaining character consistency. The images are generated in an auto-regressive way, where all previous text-image pairs are given, referred to as history text-image pairs.
1. Introduction
@@
With the rapid advancement of diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015), text-to-image generation  (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Podell et al., 2024) has achieved remarkable quality and has been widely applied to different tasks. While common text-to-image generation focuses on producing a single high-quality image from a given prompt, visual storytelling presents a more challenging setup by requiring the generation of a sequence of images that maintains consistency in characters, objects, and style across different frames. In visual storytelling, the model takes a sequence of text prompts forming a coherent storyline and generates the corresponding images. Based on whether initial reference images are provided, the task can be further categorized into two settings: story continuation, where an initial reference image is given, and story generation, where no reference image is given.
@@
Recent works have explored visual storytelling employing diffusion models in an auto-regressive manner (Pan et al., 2024; Rahman et al., 2023; Feng et al., 2023; Liu et al., 2024; Song et al., 2024; Yang et al., 2024), where, in addition to the current text prompt, all previous text-image pairs, referred to as history text-image pairs, are provided as additional conditions for generating the next image. However, one limitation of these approaches is that not all history contributes equally to the generation process. To address this, several works (Feng et al., 2023; Song et al., 2024; Liu et al., 2024) propose different methods to reweight the history text-image pairs. Despite their effectiveness, previous auto-regressive methods share another common limitation. The inclusion of history text and images as additional conditioning signals necessitates modifications to the diffusion model, resulting in the need for extensive training to adapt the diffusion model to these changes. In contrast, our approach processes history context using a novel multi-modal fusion setup, in addition to a lightweight adapter architecture—an efficient strategy demonstrated by IP-Adapter (Ye et al., 2023)—allowing seamless integration of history features without modifying the diffusion model.
@@
Another line of research focuses on subject-consistent generation, which aims to maintain visual consistency across images in a training-free, plug-and-play manner, such as StoryDiffusion (Zhou et al., 2024b), ConsiStory (Tewel et al., 2024), and OnePromptOneStory (Liu et al., 2025). Unlike auto-regressive methods, which rely on long narrative-style text descriptions, these approaches typically follow a structured prompt format of “character description + activity.” Although these methods can ensure character consistency, they have notable limitations: (1) They do not support reference images, meaning that they cannot generate specific, predefined characters. (2) To enforce consistency, the same character prompt is used in all frames, limiting the model’s ability to adapt to varying narrative descriptions or multiple character stories, thus constraining the flexibility of storytelling.
@@
A key characteristic of the visual storytelling task is the presence of history text-image pairs. These history texts and images contain rich contextual information that can effectively guide the generation of subsequent images. For example, prior images and text may include overlapping characters, objects, or implicit references that contribute to maintaining continuity and coherence throughout the story. Furthermore, history text-image pairs naturally align, as the text provides a description of the visual elements present in the images.
@@
As the storyline progresses, another challenge arises: the growing length of history increases computational and memory demands. Moreover, not all history pairs remain equally relevant, and earlier entries may introduce noise or dilute the focus of generation (Liu et al., 2024). To address this, we further introduce a salient history selection strategy that dynamically identifies the most informative history pair during inference. This improves generation quality by ensuring the model conditions only on the most relevant context.
To better evaluate how well the generated images reflect the narrative, we adopt TIFA (Hu et al., 2023), a recently proposed metric that uses vision language models to verify whether specific concepts from the text appear in the image. Unlike CLIP-based similarity scores that primarily capture high-level semantic alignment, TIFA offers more targeted and interpretable assessments of text-image alignment by focusing on explicit concept grounding.
@@
  3. (3)
    We introduce a salient history selection strategy, which dynamically selects the most informative history text-image pair during inference, improving the conditioning quality.
  4. (4)
@@
2. Related Works
2.1. Visual Story Telling
Visual storytelling involves generating a sequence of images based on narrative text prompts to create a consistent and coherent story. Several recent works have explored an auto-regressive generation process, first introduced by AR-LDM (Pan et al., 2024) where, in addition to the current text prompt, all history text-image pairs serve as additional conditions for generating the next image. Building upon this approach, Make-A-Story (Rahman et al., 2023) proposes a visual memory-attention module to enhance consistency across frames.
However, not all historical information should necessarily contribute equally to the generation process. To address this limitation, ACM-VSG (Feng et al., 2023) introduces an adaptive encoder that evaluates the semantic relationship between the current caption and history text-image pairs. It further selects the most relevant history image as an additional guidance signal during sampling to improve global consistency. Causal-Story (Song et al., 2024) takes a different approach by incorporating a local causal attention mechanism that quantifies the causal relationships between history and current text-image pairs, assigning appropriate weights to past information. StoryGen (Liu et al., 2024) enhances temporal awareness by introducing different noise levels to history images, enabling the model to better distinguish temporal order. To further improve the robustness of history features, StoryGPT-V (Shen and Elhoseiny, 2025) leverages Large Language Models (LLMs) for their strong reasoning capabilities, helping the model navigate ambiguous references. A character-aware latent diffusion model is then aligned with LLM outputs to generate character-consistent stories.
There are other methods (Zhu and Tang, 2023; Gong et al., 2023; Zhou et al., 2024a; Jeong et al., 2023; Tao et al., 2024; Zheng and Fu, 2024; Ma et al., 2024) that aim to improve story generation in different ways by introducing extra control, such as layout, stretch, identity embedding, etc. More recently, MM-Interleaved (Tian et al., 2024) and SEED-Story (Yang et al., 2024) employ a Multimodal Large Language Model (MLLM) to generate image features conditioned on history text-image pairs and the current text. These features are then used as conditioning inputs to an SDXL diffusion model for image generation. While this approach leverages the comprehension and generation ability of the MLLM to improve consistency, it significantly increases computational costs due to the need for fine-tuning both the MLLM and the diffusion model.
@@ 2.1. Visual Story Telling
Figure 2. Overview of our proposed ViSTA model. (a) To generate current image $I_{k+1}$, ViSTA takes the current text prompt $P_{k+1}$, history prompt $P_{k}$, and history image $I_{k}$ as input. We propose a multi-modal history fusion model to extract and integrate the historical information and output a fusion feature $c^{F}$. We utilize a lightweight history adapter to condition on the fusion feature, avoiding modifying the diffusion models’ architecture or full fine-tuning of the models; only the fusion model and adapter need to be trained. (b) Our proposed multi-modal history fusion model consists of $d$ blocks of history cross-attention and feedforward layers. In the history cross-attention, we use the current prompt embedding $c^{P}_{k+1}$ as the query, and the concatenation of history prompt $c^{P}_{k}$ and image embedding $c^{I}_{k}$ as the key and value. (c) Salient history selection strategy during inference. Instead of using all history prompts $P_{1,...,k}$ and images $I_{1,...,k}$ as references for the current generation, we select the most salient history based on the attention score $s$. In this way, we make the model focus on the most informative history.
2.2. Subject Consistent Generation
@@
StoryDiffusion (Zhou et al., 2024b) proposes Consistent Self-Attention, a training-free and hot-pluggable module that incorporates sampled reference tokens from the images generated within the same batch, to improve consistency. OneActor (Wang et al., 2024) derives a cluster-based score function and proposes a cluster-conditioned model, which incorporates posterior samples to guide the denoising towards the target cluster in a one-shot tuning manner. DreamStory (He et al., 2024) proposes a novel training-free multi-subject consistent model that leverages the created character portraits of the subjects and their corresponding textual information. The Chosen One (Avrahami et al., 2024) proposes iteratively generating multiple images of the character, grouping their semantic embeddings to identify the most cohesive set, and extracting the character features to generate consistent story images. OnePromptOneStory (Liu et al., 2025) proposes concatenating all prompts as a single extended prompt to maintain a consistent representation of the identity. The authors then apply Singular-Value Reweighting to refine frame descriptions and Identity-Preserving Cross-Attention to strengthen consistency at the attention level.
Unlike auto-regressive methods that rely on long narrative-style text descriptions to guide image generation, subject-consistent generation approaches typically adopt a more structured prompt format, such as “character description + activity”. This format explicitly separates the visual identity of a character from the specific actions or scenes, making it easier for the model to preserve consistency. Repetition of the same character description across different prompts, the model reinforces key visual attributes throughout the sequence. However, while effective in preserving visual coherence, these methods come with several notable limitations. (1) Lack of reference image support: These models do not incorporate reference images, meaning that they cannot generate specific predefined characters based on external visual inputs. Instead, they rely solely on textual descriptions, making it difficult to recreate a particular character with fine-grained control over their appearance. (2) Limited adaptability to narrative variations: To enforce consistency, the same character prompt must be used across all frames, which restricts the model’s ability to adapt dynamically to evolving narrative descriptions. This limitation is also problematic for complex storylines that involve multiple characters.
@@ 3. Method
3.1. Problem Setup
In this work, we focus on the visual storytelling task, where a sequence of images is generated based on narrative text captions. Given a sequence of text prompts that form a storyline, $P_{1},P_{2},\dots,P_{K}$, our goal is to generate a corresponding sequence of images, $I_{1},I_{2},\dots,I_{K}$. Following (Pan et al., 2024; Liu et al., 2024), we adopt an auto-regressive generation approach, which means that at each image generation step, all previously generated text-image pairs, along with the current text caption, are used as input. We define the $k$-th generation step as:
@@ 3.1. Problem Setup
where $k\in\{1,...,K-1\}$. Note that we focus on the story continuation setup, i.e., at the $k=1$ step, the first real image $I_{1}$ along with its caption $P_{1}$ is provided as the initial history to generate $I_{2}$.
3.2. Multi-modal History Fusion Model
@@
(1) $c^{F}=\text{Attention}(Q^{F},K^{F},V^{F})=\text{Softmax}\left(\frac{Q^{F}(K^{F})^{T}}{\sqrt{d}}\right)V^{F},$
where the query, key, and value matrices are obtained as follows:
(2) $Q^{F}=c^{P}_{k+1}W_{q}^{F},$
(3) $K^{F}=(c^{P}_{k}\|c^{I}_{k})W_{k}^{F},$
(4) $V^{F}=(c^{P}_{k}||c^{I}_{k})W_{v}^{F}.$
@@
The output fusion feature $c^{F}$ is designed to incorporate relevant history features based on the current text prompt $P_{k+1}$. This process enables the fusion feature $c^{F}$ to selectively integrate the information from the history text-image pair that is most relevant to the current prompt. In this way, it can effectively capture contextual dependencies, ensuring that historical visual and textual features are aligned with the current text, thereby enhancing consistency in the generated output.
3.3. History Adapter
Lightweight adapters designed for specific tasks are used along with the base generative models in T2I generation to enhance the generation quality for the task. One such method is IP-Adapter (Ye et al., 2023) that has demonstrated strong performance in generating consistent images conditioned on a reference image, using only lightweight trainable parameters. Inspired by this, we propose a history adapter that can be seamlessly integrated into a frozen diffusion model. The history adapter leverages the fusion feature $c^{F}$, which encodes relevant historical context, as an additional conditioning signal to guide the story image generation process.
To incorporate these features effectively, we introduce new cross-attention layers into the original U-Net (Ronneberger et al., 2015) architecture. These layers allow the image latents from the diffusion process to attend to the history fusion feature $c^{F}$. Specifically, we modify the standard cross-attention operation so that the U-Net’s intermediate latent representations serve as queries, while the history fusion feature $c^{F}$ provide the keys and values:
(5) $Z^{c}=\text{Attention}(Q^{c},K^{c},V^{c})=\text{Softmax}\left(\frac{Q^{c}(K^{c})^{T}}{\sqrt{d}}\right)V^{c}.$
@@ 3.3. History Adapter
To preserve the original behavior of the pretrained diffusion model while integrating historical context, we compute the final cross-attention output as a weighted combination of the original U-Net cross-attention output $Z$ and the history-guided output $Z^{c}$:
(6) $Z^{\prime}=Z+\lambda\cdot Z^{c},$
where $\lambda$ is a scaling factor that controls the influence of the history adapter.
@@
(7) $\mathcal{L}=\mathbb{E}_{x_{0},\epsilon,c^{P},c^{F},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},c^{P},c^{F},t)\right\|^{2}.$
Figure 3. TIFA evaluation example for the StorySalon test dataset. Given the caption, the following question-answer pairs are generated by a language model. Both binary Yes/No questions and multiple choices questions can be generated. Then a VQA model UnifiedQA evaluates the generated image based on the question-answer pairs. The final TIFA score for the generated image is the average on all questions.
3.4. Salient History Selection
@@
(8) $s_{i}=\text{Softmax}\left(\frac{Q^{F}_{i}(K^{F}_{i})^{T}}{\sqrt{d}}\right).$
@@
3.5. Text-Image Alignment Evaluation with TIFA
In visual storytelling, narrative text prompts often describe specific objects, actions, and scenes that need to be accurately reflected in the generated images. Traditional evaluation metrics like CLIP similarity focus on overall semantic similarity but may overlook whether these specific elements are visually grounded. To better assess this alignment, we adapt TIFA (Hu et al., 2023), a vision-language evaluation framework that measures how well an image supports the information described in its associated text.
As illustrated in Figure 3, TIFA uses a language model to automatically generate question-answer pairs from a given text prompt. These questions are designed to cover important content, including Yes / No questions (e.g., “Is there a baby elephant in the back of the truck?”) and multiple choice questions (e.g., “What type of vehicle are they traveling in? A. Truck, B. Car, C. Van, D. Bus”). A Visual Question Answering (VQA) model is then used to answer these questions based solely on the generated image. The TIFA score is calculated as the average accuracy across all questions, reflecting how well the image supports the textual content.
In our work, we apply TIFA to evaluate whether the generated images faithfully represents the corresponding narrative prompt. To adapt TIFA to the storytelling setting, we treat each text and generated image pair independently and compute the score at the frame level. This provides a more targeted and interpretable evaluation of how well the generated visuals align with the story captions.

### v3

@@ Abstract
Figure 1: ViSTA sample visual storytelling. Given a sequence of narrative story prompts and the first text-image pair as initial references, we generate high-quality story images aligned with their corresponding prompts, maintaining character consistency. The images are generated in an auto-regressive way, where all previous text-image pairs are given, referred to as history text-image pairs.
1 Introduction
@@
With the rapid advancement of diffusion models [8, 28], text-to-image generation  [18, 23, 26, 24, 20] has achieved remarkable quality and has been widely applied to different tasks. While common text-to-image generation focuses on producing a single high-quality image from a given prompt, visual storytelling presents a more challenging setup by requiring the generation of a sequence of images that maintains consistency in characters, objects, and style across different frames. In visual storytelling, the model takes a sequence of text prompts forming a coherent storyline and generates the corresponding images. Based on whether initial reference images are provided, the task can be further categorized into two settings: story continuation, where an initial reference image is given, and story generation, where no reference image is given.
@@
Recent works have explored visual storytelling employing diffusion models in an auto-regressive manner [19, 22, 4, 13, 30, 35], where, in addition to the current text prompt, all previous text-image pairs, referred to as history text-image pairs, are provided as additional conditions for generating the next image. However, one limitation of these approaches is that not all history contributes equally to the generation process. To address this, several works [4, 30, 13] propose different methods to reweight the history text-image pairs. Despite their effectiveness, previous auto-regressive methods share another common limitation. The inclusion of history text and images as additional conditioning signals necessitates modifications to the diffusion model, resulting in the need for extensive training to adapt the diffusion model to these changes. In contrast, our approach processes history context using a novel multi-modal fusion setup, in addition to a lightweight adapter architecture—an efficient strategy demonstrated by IP-Adapter [36]—allowing seamless integration of history features without modifying the diffusion model.
@@
Another line of research focuses on subject-consistent generation, which aims to maintain visual consistency across images in a training-free, plug-and-play manner, such as StoryDiffusion [38], ConsiStory [32], and OnePromptOneStory [14]. Unlike auto-regressive methods, which rely on long narrative-style text descriptions, these approaches typically follow a structured prompt format of “character description + activity.” Although these methods can ensure character consistency, they have notable limitations: (1) They do not support reference images, meaning that they cannot generate specific, predefined characters. (2) To enforce consistency, the same character prompt is used in all frames, limiting the model’s ability to adapt to varying narrative descriptions or multiple character stories, thus constraining the flexibility of storytelling.
@@
A key characteristic of the visual storytelling task is the presence of history text-image pairs. These history texts and images contain rich contextual information that can effectively guide the generation of subsequent images. For example, prior images and text may include overlapping characters, objects, or implicit references that contribute to maintaining consistency throughout the story. Furthermore, history text-image pairs naturally align, as the text provides a description of the visual elements present in the images.
@@
As the storyline progresses, another challenge arises: the growing length of history increases computational and memory demands. Moreover, not all history pairs remain equally relevant, and earlier entries may introduce noise or dilute the focus of generation [13]. To address this, we further introduce a salient history selection strategy that dynamically identifies the most informative history pair during inference. This improves generation quality by ensuring the model conditions only on the most relevant context.
To better evaluate how well the generated images reflect the narrative, we adopt TIFA [9], a recently proposed metric that uses vision language models to verify whether specific concepts from the text appear in the image. Unlike CLIP-based similarity scores that primarily capture high-level semantic alignment, TIFA offers more targeted and interpretable assessments of text-image alignment by focusing on explicit concept grounding.
@@
  3. 3.
    We introduce a salient history selection strategy, which dynamically selects the most informative history during inference, improving the conditioning quality.
  4. 4.
@@
2 Related Works
2.1 Visual Story Telling
Visual storytelling involves generating a sequence of images based on narrative text prompts to create a consistent and coherent story. Several recent works have explored an auto-regressive generation process, first introduced by AR-LDM [19] where, in addition to the current text prompt, all history text-image pairs serve as additional conditions for generating the next image. Building upon this approach, Make-A-Story [22] proposes a visual memory-attention module to enhance consistency across frames.
However, not all historical information should necessarily contribute equally to the generation process. To address this limitation, ACM-VSG [4] introduces an adaptive encoder that evaluates the semantic relationship between the current caption and history text-image pairs. It further selects the most relevant history image as an additional guidance signal during sampling to improve global consistency. Causal-Story [30] takes a different approach by incorporating a local causal attention mechanism that quantifies the causal relationships between history and current text-image pairs, assigning appropriate weights to past information. StoryGen [13] enhances temporal awareness by introducing different noise levels to history images, enabling the model to better distinguish temporal order. To further improve the robustness of history features, StoryGPT-V [27] leverages LLMs for their strong reasoning capabilities, helping the model navigate ambiguous references. A character-aware latent diffusion model is then aligned with LLM outputs to generate character-consistent stories.
There are other methods [40, 5, 39, 10, 31, 37, 16] that aim to improve story generation in different ways by introducing extra control, such as layout, stretch, identity embedding, etc. More recently, MM-Interleaved [33] and SEED-Story [35] employ a Multimodal Large Language Model (MLLM) to generate image features based on history text-image pairs and the current text, which are then used as conditions to guide the image generation for diffusion models. While this approach leverages the comprehension and generation ability of the MLLM to improve consistency, it significantly increases computational costs due to the need for fine-tuning both the MLLM and the diffusion model.
@@ 2.1 Visual Story Telling
Figure 2: Overview of our proposed ViSTA model. (a) To generate the current image $I_{k+1}$, ViSTA takes text prompt $P_{k+1}$, history prompt $P_{k}$, and history image $I_{k}$ as input. We propose a multi-modal history fusion model to extract and integrate the history information and output a fusion feature $c^{F}$. We utilize a lightweight history adapter to condition on the fusion feature, avoiding modifying the diffusion models’ architecture or full fine-tuning of the models; only the fusion model and adapter need to be trained. (b) Our proposed multi-modal history fusion model consists of $d$ blocks of history cross-attention and feedforward layers. In the history cross-attention, we use the current prompt embedding $c^{P}_{k+1}$ as the query, and the concatenation of history prompt $c^{P}_{k}$ and image embedding $c^{I}_{k}$ as the key and value. (c) Salient history selection strategy during inference. Instead of using all history prompts $P_{1,...,k}$ and images $I_{1,...,k}$ as references, we select the most salient history based on the attention score $s$. In this way, we make the model focus on the most informative history.
2.2 Subject Consistent Generation
@@
The Chosen One [1] and OneActor [34] propose cluster-based test-time tuning methods to optimize the character embeddings. ConsiStory [32] proposes a subject-driven shared attention block and a correspondence-based feature injection mechanism to promote consistent appearance of subjects across frames. StoryDiffusion [38] proposes Consistent Self-Attention that incorporates sampled reference tokens from the images generated within the same batch, to improve consistency. DreamStory [6] proposes a novel multi-subject consistent model that leverages the created character portraits of the subjects and their corresponding textual information. OnePromptOneStory [14] proposes concatenating all prompts as a single extended prompt and refining frame descriptions with Singular-Value Reweighting. They also propose an Identity-Preserving Cross-Attention to strengthen character consistency.
Subject-consistent generation methods typically adopt a more structured prompt format, such as “character description + activity”. This format explicitly separates the character identity from the specific actions or scenes, making it easier for the model to preserve consistency. Repetition of the same character description across different prompts, the model reinforces key visual attributes throughout the sequence. However, while effective in preserving visual coherence, these methods come with several notable limitations. (1) Lack of reference image support: they cannot generate specific predefined characters based on external visual inputs. Instead, they rely solely on textual descriptions, making it difficult to recreate a particular character with fine-grained control. (2) To enforce consistency, the same character prompt must be used across all frames, which restricts the model’s ability to adapt dynamically to evolving narrative descriptions. This limitation is also problematic for complex storylines that involve multiple characters.
@@ 3 Method
3.1 Problem Setup
In this work, we focus on the visual storytelling task, where a sequence of images is generated based on narrative text captions. Given a sequence of text prompts that form a storyline, $P_{1},P_{2},\dots,P_{K}$, our goal is to generate a corresponding sequence of images, $I_{1},I_{2},\dots,I_{K}$. Following [19, 13], we adopt an auto-regressive generation approach, i.e., at each image generation step, all previously generated text-image pairs, along with the current text caption, are used as input. We define the $k+1$-th generation step as:
@@ 3.1 Problem Setup
where $k\in\{1,...,K\}$. Note that we focus on the story continuation setup, i.e., the first real image $I_{1}$ along with its caption $P_{1}$ is provided as the initial history to generate $I_{2}$.
3.2 Multi-modal History Fusion Model
@@
$c^{F}=\text{Attn}(Q^{F},K^{F},V^{F})=\text{Softmax}\left(\frac{Q^{F}(K^{F})^{T}}{\sqrt{d}}\right)V^{F},$ (1)
where the query, key, and value are obtained as follows:
$Q^{F}=c^{P}_{k+1}W_{q}^{F},$ (2)
$K^{F}=(c^{P}_{k}\|c^{I}_{k})W_{k}^{F},$ (3)
$V^{F}=(c^{P}_{k}||c^{I}_{k})W_{v}^{F}.$ (4)
@@
The output fusion feature $c^{F}$ is designed to incorporate relevant history features based on the current text prompt $P_{k+1}$. This process enables the fusion feature $c^{F}$ to selectively integrate the information from the history text-image pair that is most relevant to the current prompt. In this way, it can effectively capture contextual dependencies, ensuring that historical visual and textual features are aligned with the current text, thereby enhancing the consistency.
3.3 History Adapter
Lightweight adapters designed for specific tasks are used along with the base generative models in T2I generation to enhance the generation quality for the task. One such method is IP-Adapter [36] that has demonstrated strong performance in generating consistent images conditioned on a reference image, using only lightweight trainable parameters. Inspired by this, we propose a history adapter that can be seamlessly integrated into a frozen diffusion model. The history adapter leverages the fusion feature $c^{F}$, which encodes relevant historical context, as an additional conditioning signal to guide the story image generation process.
To incorporate these features effectively, we introduce new cross-attention layers into the original U-Net [25] architecture. These layers allow the image latents from the diffusion process to attend to the history fusion feature $c^{F}$. Specifically, we modify the standard cross-attention operation so that the U-Net’s intermediate latent representations serve as queries, while the history fusion feature $c^{F}$ provide the keys and values:
$Z^{c}=\text{Attn}(Q^{c},K^{c},V^{c})=\text{Softmax}\left(\frac{Q^{c}(K^{c})^{T}}{\sqrt{d}}\right)V^{c}.$ (5)
@@ 3.3 History Adapter
To preserve the original behavior of the pretrained diffusion model while integrating historical context, we compute the final cross-attention output as a weighted combination of the original U-Net cross-attention output $Z$ and the history-guided output $Z^{c}$, where the scaling factor $\lambda$ controls the influence of the history adapter:
$Z^{\prime}=Z+\lambda\cdot Z^{c},$ (6)
@@
$\mathcal{L}=\mathbb{E}_{x_{0},\epsilon,c^{P},c^{F},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},c^{P},c^{F},t)\right\|^{2}.$ (7)
3.4 Salient History Selection
@@
$s_{i}=\text{Softmax}\left(\frac{Q^{F}_{i}(K^{F}_{i})^{T}}{\sqrt{d}}\right).$ (8)
@@
Figure 3: TIFA evaluation example. Given the caption, the following question-answer pairs are generated by a language model. Then a VQA model UnifiedQA evaluates the generated image based on the question-answer pairs. The final TIFA score for the generated image is the average on all questions.
3.5 Text-Image Alignment Evaluation with TIFA
In visual storytelling, narrative text prompts often describe specific objects, actions, and scenes that need to be accurately reflected in the generated images. Traditional evaluation metrics like CLIP similarity focus on overall semantic similarity but may overlook whether these specific elements are visually grounded. To better assess this alignment, we adapt TIFA [9], a vision-language evaluation framework that measures how well an image supports the information described in its associated text.
We follow the question generation pipeline proposed in the original TIFA paper [9], as shown in Figure 3. Specifically, a language model is used to automatically generate question-answer pairs from a given story caption, guided by in-context examples. These questions are designed to cover important content, including Yes / No questions (e.g., “Is there a baby elephant in the back of the truck?”) and multiple choice questions (e.g., “What type of vehicle are they traveling in? A. Truck, B. Car, C. Van, D. Bus”). These questions target specific entities and activities mentioned in the caption. Then, a Visual Question Answering (VQA) model is used to answer these questions based solely on the generated image. The TIFA score is calculated as the average accuracy across all questions, reflecting how well the image supports the textual content.
In our work, we apply TIFA to evaluate whether the generated images faithfully represents the corresponding narrative prompt. To adapt TIFA to the storytelling setting, we treat each text and generated image pair independently and compute the score at the frame level. This provides a more targeted and interpretable evaluation of how well the generated visuals align with the story captions, especially when the story captions are long and rich in narrative structure.